A LIQUID TAKES 10 MINUTE TO COOL FROM 70C TO 50C HOW MUCH TIME WILL IT TAKE TO COOL FROM 60C TO 80C THE TEMPRATURE OF SURROUNDING IS 20C

Grade:12

1 Answers

vikas askiitian expert
13 years ago

a simple method for these types of problem is average temperature method ...

rate of cooling = dT/dt = k [(Ti+Tf)/2 - To]

To is temperature of surrounding ...

 

liquid takes 10minutes to cool from 70 to 50 so

rate = 70-50/10 = k [ (70+50)/2 -20 ]

 2 = k[ 60-20]

k = 1/20              ...................1

 

let time taken by liquid to cool from 80 to 60 is t then

rate = dT/t = 20/t = k [ (80+60)/2 - 20]

 20/t = 50k

  t = 2/5k

k = 1/20 from eq 2 so

t = 8 mins

this is the required time , approve my ans if u like